OpenStudy (anonymous):

Will FAN AND GIVE MEDAL 5+2(4X-8)=-X-5(3X-2)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

The answer will be: X=7/8

OpenStudy (anonymous):

explanation please @damneny

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

Expand: 5+8X−16=−X−15X+10 2) Simplify 5+8X−16 to 8X−11 8X−11=−X−15X+10 3) Simplify −X−15X+10 to −16X+10 8X−11=−16X+10 4) Add 11 to both sides 8X=−16X+10+11 5) Simplify −16X+10+11 to −16X+21 8X=−16X+21 6) Add 16X to both sides 8X+16X=21 7) Simplify 8X+16X to 24X 24X=21 8) Divide both sides by 24 X=21/24 9) Simplify 21/24 to 7/8 X=7/8

OpenStudy (anonymous):

is this from http://www.mathpapa.com/algebra-calculator.html by any chance?

OpenStudy (anonymous):

i like your thinking but don't get it

OpenStudy (anonymous):

Ok, one sec.

OpenStudy (anonymous):

Expand this: 5+2(4X-8)=-X-5(3X-2), and you'll get 5+8x-16 = -x-15x+10

OpenStudy (anonymous):

Shift the +5 and -16 to the other side, so then the +5 will become -5 and the -16 will change to +16. So: 8x=-16x+10-5+16

OpenStudy (anonymous):

Shift the -16x to the left side and it'll become positive 16x, so: 8x+16x=10-5+16

OpenStudy (anonymous):

So, we will get after solving these: 24x=21

OpenStudy (anonymous):

Then finally, we will get x= 21/24. They are divisible by 3 so, divide them by 3. The final answer will be x=7/8. I hope it helped :/